amanda has 20 coins worth a total of $1.35. the coins are all nickels and dimes. how many coins of each does she have ?
Answer by mananth(16949)

5x-10x+ 200=135
-5x+200 =135
-5x=135-200
-5x=-65
/-5
x=13 nickels
balance 7 dimes